1. The following data represent the age of first arrest for a sample of 50 boys.

X (age)

f

cf

%

c%

9

1

1

2%

2%

10

9

10

18%

20%

11

10

20

20%

40%

12

5

25

10%

50%

13

7

32

14%

64%

14

4

36

8%

72%

15

7

43

14%

86%

16

5

48

10%

96%

17

1

49

2%

98%

19

1

50

2%

100%

Σ

50

100

a. What is the mode, median and mean of this data? Which is the best measure of central tendency? Explain. Without graphing the data, what can you assume the distribution will look like? What makes you think this?

B. You take a sample of 50 police officers from large urban police departments in Maryland. The table below provides information on your sample of 50 officers and the number of civilian complaints filed against those officers over the past 10 years. Use this ungrouped frequency distribution to answer the questions below.

# Complaints

f

cf

%

c%

0

19

19

38

38

1

7

26

14

52

2

8

34

16

68

3

6

40

12

80

5

4

44

8

88

6

2

46

4

92

7

2

48

4

96

9

1

49

2

98

13

1

50

2

100

Σ

50

100

Calculate the Mean, Median, and Mode of the data.
Answer #1

As shown above we can find the mean median mode for below data as well

Mode=0 as it comes max times

Median=1 (As both 25th and 26th observations are 1)

Mean= Sigma (Complaint*f)/50 = 109/50 =2.18

Hope the above answer has helped you in understanding the problem. Please upvote the ans if it has really helped you. Good Luck!!